Solve the inequalities. Give your answer in interval notation, and indicate the answer


geometrically on the real-number line.


a. t + 6 ≤ 2 + 3t


b. 3(2 – 3x) > 4(1 – 4x)



1
Expert's answer
2022-05-24T23:15:03-0400

a.



"t + 6\le 2 + 3t""3t-t\ge6-2""2t\ge4""t\ge 2"

"t\in [2, \infin)"


b.



"3(2 \u2013 3x) > 4(1 \u2013 4x)""6-9x>4-16x""6-9x+16x-6>4-16x+16x-6""7x>-2""x>-\dfrac{2}{7}""x\in (-\dfrac{2}{7}, \infin)"